can someon help me solve the following rational equation: x+1 over 1 = 3 over 2
Answer by stanbon(75887)

solve the following rational equation: x+1 over 1 = 3 over 2
-----
(x+1)/1 = 3/2
---
Cross-multiply to get:
2(x+1) = 3
2x+2 = 3
2x = 1
x = 1/2
--------------
Cheers,
Stan H.
===============